OpenStudy (anonymous):

what are the first five terms of the following sequence? an = 2 - n

OpenStudy (solomonzelman):

a1=2-1 a2=2-2 a3=2-3 so the difference is , d=-1. I already posted a2 and a3, can you at least tell me a4 ?